### v1

# CASA: Class-Agnostic Shared Attributes in Vision-Language Models for Efficient Incremental Object Detection

Abstract
Incremental object detection (IOD) is challenged by background shift, where background categories in sequential data may include previously learned or future classes. Inspired by the vision-language foundation models such as CLIP, these models capture shared attributes from extensive image-text paired data during pre-training. We propose a novel method utilizing attributes in vision-language foundation models for incremental object detection. Our method constructs a Class-Agnostic Shared Attribute base (CASA) to capture common semantic information among incremental classes. Specifically, we utilize large language models to generate candidate textual attributes and select the most relevant ones based on current training data, recording their significance in an attribute assignment matrix. For subsequent tasks, we freeze the retained attributes and continue selecting from the remaining candidates while updating the attribute assignment matrix accordingly. Furthermore, we employ OWL-ViT as our baseline, preserving the original parameters of the pre-trained foundation model. Our method adds only 0.7% to parameter storage through parameter-efficient fine-tuning to significantly enhance the scalability and adaptability of IOD. Extensive two-phase and multi-phase experiments on the COCO dataset demonstrate the state-of-the-art performance of our proposed method.
1 Introduction
State-of-the-art object detection models [10, 36, 35, 13, 3, 26] have achieved remarkable success in recognizing and localizing objects within a fixed set of categories. These models operate under a critical limitation: they assume a static training environment where all object categories are known and labelled beforehand. In real-world scenarios, as new object categories emerge over time, the need for continuous updates to these models becomes crucial—a process known as incremental object detection (IOD) [37, 9, 24]. However, IOD also suffers significant challenges, particularly in managing background drift [4]. Background drift occurs when objects belonging to previous or future tasks in IOD are not annotated in the current task and are instead assigned to the background class. This can cause serious misclassification issues [5, 24], as the model may later confuse these background objects with newly introduced categories or fail to recognize them altogether. This highlights a key challenge in incremental learning: how to maintain the ability to generalize across tasks without comprehensive annotations.
In traditional object detection [11, 22, 53], accurate classification relies on extracting and utilizing shared visual features, i.e., shape, texture, and color, across categories. These fine-grained semantic details are crucial for distinguishing objects from the background, particularly as new categories appear. Research [31, 40, 2] in transfer learning and domain adaptation highlights that effective generalization depends on the ability to leverage these common attributes. However, in IOD, the evolving background class with each new task complicates the traditional models struggle to consistently capture and apply these shared features.
Recent advances in vision-language models [32, 59, 58, 46, 43, 50, 51, 48, 49] offer promising solutions to these challenges. By integrating visual and textual data, these models provide a richer contextual understanding of both objects and their backgrounds. This cross-modal approach enhances the ability to retain and align shared semantic information across tasks, thereby mitigating the effects of background drift and improving the robustness of incremental object detection. The inability to maintain a coherent representation of these semantic relationships across tasks is thus a fundamental problem in IOD, compromising its overall performance as it encounters new categories.
In response to these challenges, we propose a novel approach that leverages vision-language foundation models to address the background drift problem in IOD. Our method named Class-Agnostic Shared Attributes (CASA), a framework designed to capture and utilize common semantic information across incremental classes. By employing large language models, we generate candidate textual attributes relevant to the object categories in each detection task, curate these attributes based on their relevance to the current training data, and record their importance in an attribute assignment matrix. For subsequent tasks, we freeze the retained attributes while continuing to select and update relevant attributes, ensuring that the model adapts incrementally without losing previously learned knowledge. Moreover, we utilize OWL-ViT [30] as the baseline, preserving the original parameters of the pre-trained foundation model. Our method achieves only a minimal increase in parameter storage (0.7%) through parameter-efficient fine-tuning, significantly enhancing the scalability and adaptability of IOD. Extensive experiments on the COCO dataset demonstrate the effectiveness of our approach, achieving state-of-the-art performance across both two-phase and multi-phase incremental learning scenarios.
In summary, our contributions are as follows:
  • We propose the Class-Agnostic Shared Attributes (CASA) for leveraging common semantic information across categories in incremental object detection, enhancing the ability to handle background drift.
  • Our method utilizes a frozen vision-language foundation model with parameter-efficient fine-tuning that only increases parameter storage by 0.7%, significantly improving the scalability and adaptability of IOD.
  • Extensive two-phase and multi-phase experiments on the COCO dataset demonstrate the effectiveness and efficiency of our approach, achieving state-of-the-art performance in incremental object detection.
2 Related Works
2.1 Vision-Language Models
Vision-language models have emerged as powerful tools for understanding and integrating visual and textual data, enabling more comprehensive and context-aware representations of objects and scenes. One of the most prominent models is CLIP (Contrastive Language-Image Pretraining) [32], which leverages large-scale image-text pairs to learn a joint embedding space where visual and textual modalities are aligned. This approach has inspired subsequent research in vision-language models, such as ALIGN [14] and Florence [47], which further enhance the ability to generalize across diverse datasets and tasks. Recent advancements in vision-language models about object detection and classification have significantly improved robustness and adaptability, particularly in scenarios involving novel or evolving categories [32, 58, 60]. Among these methods, OWL-ViT [30], an open-world learning vision transformer, stands out by enabling more flexible and scalable detection and classification. FOMO [61] based on OWL-ViT utilize the attributes of known objects to recognize unknown objects in the open world environment.
2.2 Incremental Learning Based on Pre-trained Models
Traditional incremental learning methods [6, 19, 21] usually start from a model trained from scratch. However, with the emergence of a variety of foundation models, leveraging their strong generalization to downstream tasks, incremental learning with pre-trained models has become an increasingly popular topic. It can be divided into three strategies: Prompt-based, Representation-based, and Model Mixture-based methods [56]. Prompt-based methods leverage the strong generalization capabilities of pre-trained models by using prompts to perform lightweight updates without fully fine-tuning all model parameters [42, 44, 45, 38, 15, 39, 16]. Representation-based methods directly utilize the generalization capabilities of PTMs to construct classifiers without making significant adjustments to the model itself [57, 28, 1, 52, 54, 12]. Model Mixture-based methods design a set of models during the learning process and employ techniques like model merging and ensemble learning to make final predictions [55, 41, 27]. By combining different models, these methods aim to capitalize on the strengths of each model to enhance overall learning performance.
2.3 Incremental Object Detection
Incremental Object Detection (IOD) [37, 9] presents unique challenges compared to standard object detection, particularly in managing the evolving nature of the background class and ensuring that newly introduced categories do not interfere with previously learned ones. Existing incremental object detection methods can be broadly classified into three categories [63]: Knowledge Distillation (e.g., LWF [20]), Replay(e.g., iCaRL [33], ABR [24], and Regularization methods(e.g., ERD [9]). In recent years, with the popularity of the Transformer architecture, more and more methods try to use DETR as the baseline. CL-DETR [25] extends the DETR architecture to support incremental learning scenarios by leveraging transformer-based representations. Furthermore, CIOD [17] specifically focuses on maintaining detection accuracy across a growing number of classes by integrating adaptive feature extractors and regularization strategies. Additionally, the integration of vision-language models into IOD [18, 7] is a promising direction. These methods demonstrate that textual prompts can provide more semantic information for the representation of visual modality.
Figure 1: Illustration of our proposed Class-Agnostic Shared Attribute (CASA) in Vision-Language Models for Efficient Incremental Object Detection. We leverage large language models to generate candidate textual attributes $E_{a}$ and then select the most relevant ones $\hat{E}_{a}^{t}$ based on the current training data, documenting their significance in an attribute assignment matrix $A^{t}$. In subsequent tasks, we retain and freeze these selected attributes, continuing the process by choosing from the remaining candidates and appending them after $\hat{E}_{a}^{t-1}$, and updating the attribute assignment matrix accordingly.
3 Method
3.1 Problem Formulation and Overview
Incremental object detection is an extension of object detection that allows for learning new categories without forgetting old categories. Specifically, when the model learns new classes during Task $T_{t}$, it must also maintain the ability to recognize the old classes ranging from Task $T_{1}$ to Task $T_{t-1}$. However, the challenge of background shift happens when objects from previous or future tasks in IOD are not labeled in the current task and are mistakenly categorized as part of the background class. Therefore, we propose an innovative method that leverages shared attributes within vision-language foundation models to enhance incremental object detection shown in Figure 1.
3.2 Attribute and Assignment Matrix Generation
Before encoding the text information, we first generate a large amount of attribute information [29, 62] corresponding to different object categories using the large language model (i.e., Gpt-3.5). Then we apply the attribute prompt template to obtain $N$ pieces of attribute text information, such as “object which (is/has/etc) Shape is squared.” Subsequently, we encode these texts using a Text Encoder [32] to obtain the attribute embedding, denoted as $E_{a}=[\mathbf{e}^{t}_{A_{1}},\mathbf{e}^{t}_{A_{2}},...,\mathbf{e}^{t}_{A_{N}}]\in\mathbb{R}^{D\times N}$. The attribute embedding $E_{a}$ is shared by current task and future tasks. We also use the vision encoder during each task to encode image information [8]. In our vision-language model, it is used together with $E_{a}$ for our efficient incremental object detection.
It is worth mentioning that not all attribute embedding is used in the detection process, and the selected attribute embedding has specific weights for each class. Therefore, assigning a score to represent the attribute assignment for each class during the selection becomes particularly important. Therefore, we introduce an assignment matrix, where each element represents the score of an attribute embedding assigned in the detection process for a certain class. In the process of incremental object detection, we denote the assignment matrix in $T_{t}$ as $A^{t}\in\mathbb{R}^{N\times C^{1:t}}$, where $C^{1:t}$ represents the number of all categories learned until Task $T_{t}$, while $C^{t}$ represents the number of new categories in Task $T_{t}$. In each task, We first initialize this assignment matrix with arbitrary values between 0 and 1 for subsequent training.
3.3 Assignment Matrix Updating
In Task $T_{t}$, with attribute embedding $\mathbf{e}^{t}_{A}$ and visual embedding $\mathbf{e}^{v}$ [8], we can calculate the similarity vector $S^{t}$ as follows:
$S^{t}=\text{CosSim}(\mathbf{e}^{v},\mathbf{e}^{t}_{A}).$ (1)
Subsequently, we can use the similarity vector $S^{t}$ to update the assignment matrix $A^{t}$. We take the newly introduced last $C^{t}$ columns of the matrix $A^{t}$ as $A^{t-1:t}$. Then, the class probability $p_{cls}$ corresponding to $C^{t}$ classes introduced in Task $T_{t}$ can be computed using the following equation [23]:
$p_{cls}=\text{Sigmoid}(A^{t-1:t}S^{t}).$ (2)
Next, we can calculate the Binary Cross-Entropy (BCE) loss between the logits of these categories and the one-hot encoded targets, $y$, to update the assignment matrix $A^{t-1:t}$. Additionally, we should add an regularization loss to enforce sparsity, resulting in a more sparse matrix $A^{t-1:t}$. The loss function $\mathcal{L}_{upd}$ can be formulated as:
$\mathcal{L}_{upd}=\mathcal{L}_{BCE}(p_{cls},y)+\lambda_{1}\sum_{i,j}|A^{t-1:t}_{t_{i,j}}|,$ (3)
where $\lambda_{1}$ is a tunable hyperparameter, which is set to 0.01 fixed in our work.
At this stage, we have obtained an assignment matrix $A^{t-1:t}\in\mathbb{R}^{N\times C^{t}}$ with elements ranging between 0 and 1. Next, we proceed to filter the shared attributes. We assume $H_{a}$ represents the number of representative attributes selected for each category, consistently set to 25 in our experiments. We flatten $A^{t-1:t}$ into a one-dimensional vector with $N\times C^{t}$ elements, from which we select the top $C^{t}\times H_{a}$ elements with the highest values. The remaining elements are set to zero. Afterwards, we reshape $A^{t-1:t}$ back into a two-dimensional matrix where all elements are either 0 or 1.
For the first task, we save the assignment matrix $A^{1}$. For the following tasks, to implement incremental learning, we concatenate the $A^{t-1}$ matrix, which is saved from the previous task, with the current task’s $A^{t-1:t}$ matrix row-wise to form a new $N\times C^{1:t}$ binary matrix $A^{t}$. In Task $T_{t}$, it is essential that the assignment matrix $A^{t}$ matrix is saved for following incremental object detection.
3.4 Attribute and Assignment Matrix Selection
At this point, in Task $T_{t}$ each value in a row of the assignment matrix $A^{t}$ represents whether a particular attribute is meaningful for detecting a specific category. If the value in a row corresponding to a category is 1, it signifies that this attribute is relevant for detecting that category. Conversely, the value of 0 indicates irrelevance. We then mask the attributes according to $A^{t}$ that all values in the row of the assignment matrix $A^{t}$ are 0.
We iterate through each row of the $A^{t}$ matrix and find $\hat{N}^{t}$ rows where the elements are not all zeros. In principle, we would select $C^{1:t}\times H_{a}$ rows attribute embedding $E_{a}$ in Task $T_{t}$, but actually how much we select is far fewer than $C^{1:t}\times H_{a}$. This is because some attributes are shared among classes within Task $T_{t}$. More importantly, as incremental learning progresses, an increasing number of attributes are shared between Task $T_{t}$ and all previous tasks. We need to retain these rows and delete the rows that are all zeros. Thus, we look for the indices of the rows where the element values are not all zeros, denoted as $id^{raw}_{t}$. To achieve incremental object detection, we refer to the indices $id_{t-1}$ saved from the previous task. We then check whether each index in $id^{raw}_{t}$ appears in $id_{t-1}$. The indices that do not appear in $id_{t-1}$ are added to the end of $id_{t-1}$, forming the indices $id_{t}$ for the current task. Therefore, to achieve incremental object detection for Task $T_{t}$, we need to save $id_{t-1}$ in Task $T_{t-1}$.
Next, we use $id_{t}$ to filter and select the attribute embedding $E_{a}$ into $\hat{E}_{a}^{t}$ and assignment matrix for Task $T_{t}$, retaining only the rows corresponding to the indices recorded in $id_{t}$ to form the $\hat{E}_{a}^{t}$ and $A^{t}$. During this filtering process, we may find that some indices in $id_{t}$ also appear in $id_{t-1}$, which indicates that the attributes selected for the current task have already been selected in previous tasks. This achieves attribute sharing under class-agnostic conditions.
3.5 Adaption for Attribute Refinement
After this step, we refine the attribute embedding $\hat{E}_{a}^{t}$ based on $A^{t-1:t}$ for the current task. However, the textual attribute information generated from the text and the visual information are orthogonal in space, belonging to different domains. Therefore, before fine-tuning the textual attribute information for visual information, we adapt the attribute embedding $\hat{E}_{a}^{t}$ to eliminate barriers between visual and textual information. For each category in the current task, we take $M$ samples and calculate the visual mean embedding $\bar{E}^{v}$ for these $M$ samples:
$\bar{E}^{v}(c)=\frac{1}{M}\sum_{i=0}^{M}\mathbf{e}^{v}_{i}.$ (4)
Our adapting strategy for $\hat{E}_{a}^{t}$ aims to align the transpose of $A^{t-1:t}$ with $\hat{E}_{a}^{t}$. Additionally, to achieve incremental object detection, in Task $T_{t}$ we need to ensure that the first $Q$ rows of $\hat{E}_{a}^{t}$, $\hat{E}_{a}^{t}[\ :Q,:]$, is as consistent as possible with $\hat{E}_{a}^{t-1}$, where $Q$ equals the number of rows in $\hat{E}_{a}^{t-1}$. This can mitigate the forgetting of previously learned classes.The loss function $\mathcal{L}_{ada}$ can be formulated as:
$\mathcal{L}_{ada}=$ $\ \mathcal{L}_{MSE}\left(\bar{E}^{v}\ ,{A^{t-1:t}}^{\top}\otimes\hat{E}_{a}^{t}\right)$ (5)
$+\lambda_{2}\mathcal{L}_{MSE}\left(\hat{E}_{a}^{t}[\ :Q,:],\hat{E}_{a}^{t-1}\right),$
where $\lambda_{2}$ is also a tunable hyperparameter. In this way, we adapt the attribute embedding $\hat{E}_{a}^{t}$, eliminating barriers between visual and textual information, which can be used for future refinement.
3.6 Refinement for Inference
In Task $T_{t}$, after adapting the attribute information, we perform refining on a small scale to better apply the assignment matrix $A^{t}$ and the attribute embedding $\hat{E}_{a}^{t}$ to the subsequent inference. Similar to Equation 1, we first compute the similarity vector $\hat{S}_{t}$ between the attribute embedding $\hat{E}_{a}^{t}$ and visual embedding $\mathbf{e}^{v}$ in the current task. Next, with the assignment matrix $A^{t-1:t}$, logits corresponding to these $C_{t}$ categories can be computed using Equation 2. Unlike the previous process, at this stage, we need to keep the assignment matrix $A^{t}$ unchanged and update $\hat{E}_{a}^{t}$. We calculate the BCE loss between the probabilities $P$ of these categories and the targets $T$, which are a list of labels corresponding to the target categories in the image.
Table 1: CASA results (%) on COCO 2017 in two-phase setting 70+10. The best performance is highlighted in bold.
Scenarios Method Baseline $AP$ $FPP$ $AP_{.5}$ $FPP$ $AP_{.75}$ $FPP$
80 Joint Training OWL-ViT 42.1 non 61.8 non 47.1 non
70 Deformable DETR 43.4 non 62.8 non 47.2 non
OWL-ViT 43.76 non 63.43 non 48.37 non
70 + 10 ERD [9] UP-DETR 36.2 54.8 39.3
CL-DETR [25] UP-DETR 37.6 56.5 39.4
LwF [20] Deformable DETR 24.5 36.6 26.7
iCaRL [34] Deformable DETR 35.9 52.5 39.2
ERD [9] Deformable DETR 36.9 55.7 40.1
CL-DETR [25] Deformable DETR 40.1 57.8 43.7
VLM-PL [18] Deformable DETR 39.8 58.2 43.3
CIOD [17] Deformable DETR 40.9 1.9 59.5 2.2 44.8 1.8
CASA OWL-ViT 42.2 -0.01 61.0 -0.01 46.6 -0.01
To achieve incremental object detection, we also need to ensure that the first $Q$ rows of the current task’s $\hat{E}_{a}^{t}$, $\hat{E}_{a}^{t}[\ :Q,:]$, remains consistent with the $\hat{E}_{a}^{t-1}$ saved from the previous task. Therefore, an additional BCE loss between them is added. The loss function at this stage can be expressed as:
$\mathcal{L}_{ref}=$ $\ \mathcal{L}_{BCE}\left(P,U\right)+\lambda_{3}\mathcal{L}_{MSE}\left(\hat{E}_{a}^{t}[\ :Q,:],\hat{E}_{a}^{t-1}\right),$ (6)
where $\lambda_{3}$ is also a tunable hyperparameter. After refining the attribute embedding, the $\hat{E}_{a}^{t}$ and the $A^{t}$ can be used in the following inference stage.
In the inference phase, similar to the process described in Equation 1 and Equation 2, the image to be inferred is first encoded by the Vision Transformer Encoder to obtain visual embedding $\mathbf{e}^{v}$. The encoded vectors $\mathbf{e}^{v}$ are then multiplied by the $\hat{E}_{a}^{t}$ obtained during model training, resulting in the similarity vectors. Subsequently, they are multiplied by the assignment matrix $A^{t}$ to obtain the probability of the specific category.
4 Experiments
4.1 Dataset and Evaluation Metrics
We evaluate our method on the MS COCO 2017 dataset, which is widely used in recent incremental object detection work. Via various experimental settings, both in two-phase and multi-phase, we conduct a comprehensive comparison with other incremental object detection methods in terms of evaluation metrics $AP$, $AP_{.5}$ and $AP_{.75}$, demonstrating the superiority of our method. We also evaluate the metric called Forgetting Percentage Points ($FPP$) followed by CL-DETR [25]: $FPP=AP^{1}-AP^{t}_{old}$, where $AP^{1}$ evaluates for all classes in the first task and $AP^{t}_{old}$ evaluates for previous classes learned from the first task during current task. Moreover, through the comparison of False Positives ($FP$), we show that our method CASA effectively overcomes the background shift problem [24] in incremental object detection.
4.2 Implementation Details
Our method CASA is based on OWL-ViT, which combines a Vision Transformer (ViT) with a text encoder, allowing the model to understand both images and text prompts, designed for open-world object detection and classification. All experiments are performed using 8 NVIDIA A100 GPUs.
It is noteworthy that we do not retrain the OWL-ViT model, instead we achieve efficient fine-tuning of the pre-trained model. The parameter storage number of OWL-ViT is $4.31\times 10^{8}$, while our fine-tuned method increases the parameter storage number to $4.34\times 10^{8}$, representing an increase of less than 0.7%.
[…]
4.3 Results and Analyses
Table 2: CASA results (%) on COCO 2017 in two-phase setting 40+40. The best performance is highlighted in bold.
Scenarios Method Baseline $AP$ $AP_{.5}$ $AP_{.75}$
80 Joint OWL-ViT 42.1 61.8 47.1
40 Def-DETR 46.5 68.6 51.2
OWL-ViT 46.0 65.7 50.8
40 + 40 ERD UP-DETR 35.4 55.1 38.3
CL-DETR UP-DETR 37.0 56.2 39.1
LwF Def-DETR 23.9 41.5 25.0
iCaRL Def-DETR 33.4 52.0 36.0
ERD Def-DETR 36.0 55.2 38.7
CL-DETR Def-DETR 37.5 55.1 40.3
VLM-PL Def-DETR 41.7 59.9 44.2
CIOD Def-DETR 43.0 62.1 47.1
Ours OWL-ViT 43.2 62.5 47.2
Table 3: CASA results ($AP$/$AP_{.5}$, %) on COCO 2017 in multi-phase setting. The tasks are categorized into two distinct scenarios: 40+10+10+10+10 and 40+20+20. Our method utilizes OWL-ViT as the baseline, while others are based on deformable DETR. The best performance is highlighted in bold.
Method $\mathcal{T}_{1}$ (1-40) 40+10+10+10+10 40+20+20
$\mathcal{T}_{2}$ (40-50) $\mathcal{T}_{3}$ (50-60) $\mathcal{T}_{4}$ (60-70) $\mathcal{T}_{5}$ (70-80) $\mathcal{T}_{2}$ (40-60) $\mathcal{T}_{3}$ (60-80)
ERD 46.5 / 68.6 36.4 / 53.9 30.8 / 46.7 26.2 / 39.9 20.7 / 31.8 36.7 / 54.6 32.4 / 48.6
VLM-PL 41.7 / 59.3 38.5 / 56.4 34.7 / 53.6 31.4 / 50.8 41.7 / 60.4 39.7 / 56.5
CIOD 42.3 / 62.8 40.6 / 60.2 40.0 / 59.0 36.8 / 54.7 42.5 / 62.2 41.1 / 59.5
Ours 46.0 / 65.7 45.5 / 66.4 43.1 / 62.5 43.2 / 62.3 41.5 / 59.7 43.0 / 62.5 41.6 / 60.0
4.3.1 Two-phase setting.
In the Two-phase setting, we randomly divide the 80 classes of the COCO dataset into two experimental settings: 70+10 and 40+40. In Tab 1 and Tab 2 we compare the performance of our method, CASA, with other incremental object detection methods in terms of the metrics $AP$, $AP_{.5}$ and $AP_{.75}$ in both experimental settings. Additionally, in the 70+10 setting, we compare the difference in $FPP$ between our method and the previously best incremental object detection method, demonstrating that our method not only prevents forgetting but also achieves better performance than the previous task. As for False Positives ($FP$), in the 70+10 setting our method CASA has 31052 errors, which demonstrate an clear advantage, reducing at least 5000 errors than other methods. This indicates that our method effectively overcomes the issue of background shift compared to other incremental object detection methods. Based on the OWL-ViT model, in the first phase, the performance on 70 classes of object detection is slightly better than the previous optimal baseline Deformable DETR, while it is slightly inferior to Deformable DETR in the 40 classes setting. However, both in 70+10 and 40+40 settings, CASA consistently outperforms the state-of-the-art, better than CIOD and other incremental object detection approaches. In the 70+10 setting, our method achieves a 1.5% improvement in $AP_{.5}$ and a 1.9% improvement in $FPP$ over the current best methods, which is significantly more evident than the performance gains in the 40+40 setting. This is because our method CASA leverages class-agnostic shared attribute information for incremental object detection. The more classes learned in the first phase, the richer the preserved shared attribute information, which benefits the subsequent phases of incremental learning.
Table 4: Ablation experiments results of different components on COCO 2017 in multi-phase setting 40+10+10+10, with the best results highlighted in bold. The “—” indicates that the model lacks object detection capability from the first task, making it meaningless to calculate the FPP value in this context.
Sel Ada Ada+Sha Ref Ref+Sha 40 40+10 40+10+10 40+10+10+10 40+10+10+10+10
All All Old $FPP$ All Old $FPP$ All Old $FPP$ All Old $FPP$
$\checkmark$ 0.13 0.09 0.11 0.08 0.09 0.12 0.08 0.12 0.13
$\checkmark$ $\checkmark$ 64.61 14.02 0.18 64.43 7.62 0.16 8.31 9.58 0.15 7.47 6.70 0.15 9.43
$\checkmark$ $\checkmark$ 0.16 0.12 0.13 0.07 0.08 0.11 0.08 0.12 0.13
$\checkmark$ $\checkmark$ $\checkmark$ 65.67 14.04 0.10 65.57 7.99 0.13 13.91 9.56 0.09 7.90 6.74 0.12 9.44
$\checkmark$ $\checkmark$ 64.61 65.21 64.98 -0.22 61.11 65.43 -0.48 60.59 61.59 -0.29 57.75 60.88 -0.29
$\checkmark$ $\checkmark$ 0.15 0.12 0.15 0.15 0.15 0.17 0.16 0.15 0.16
$\checkmark$ $\checkmark$ $\checkmark$ 65.67 66.33 65.67 0.03 62.54 66.30 -0.02 62.32 62.56 0.07 59.67 62.25 0.07
4.3.2 Multi-phase setting.
In the multiple phases setting, we conducted experiments with configurations of 40+20+20 and 40+10+10+10+10 respectively. As before, the categories for each phase are randomly assigned from the COCO dataset’s 80 categories. Tab 3 records the $AP$ and $AP_{.5}$ at each step. We observe that our method, CASA, exhibits the best performance among incremental object detection methods in both multiple-phase experimental settings, even though the detection performance in the first phase is inferior to that of the method based on Deformable DETR. Notably, from the results of the 40+10+10+10 experiment, our method CASA demonstrates strong continuous learning capabilities across multiple steps, with $AP$ and $AP_{.5}$ significantly improved by 4.7% and 5%, respectively, compared to the current state-of-the-art method. This effectively indicates that our method can be applied to real-world scenarios for continuous learning.
4.4 Class-agnostic Shared Attributes
Our method, CASA, achieves incremental object detection through class-agnostic shared attributes in vision-language model. Here, we illustrate the process of increasing shared attributes using the 70+10 experimental setting as an example. In the first phase, we assume that each category is represented by 25 attribute information. In the first task, CASA only select 1314 attributes out of 2895 possible attributes for 70 categories, indicating that many attributes are shared among different categories within the first phase. In the second phase, CASA learns 10 additional categories, and in practice only 155 new attributes are added. This is partly due to some attributes being shared among these 10 categories, and more importantly, the attributes of these 10 categories are already used among the previous 70 categories, achieving efficient attribute sharing between the previous and current tasks.
Figure 2 shows the attribute scores in the two-phase setting. We observe that the first task exhibits high scores across the first five attributes, and the last five classes in the second task show high scores for some of these attributes, indicating that certain attributes are shared across different tasks. In the second task, the scores for the subsequent five attributes are notably higher. Some of these attributes had lower scores in the first task, suggesting that the second task would select a new set of shared attributes.
Figure 2: Attribute scores in two-phase setting.The first five classes belong to the initial phase, while the latter five classes are part of the second phase.
4.5 Ablation Experiments
We conduct ablation studies from two perspectives. First, we perform ablation experiments on our three main modules: updating the assignment matrix $A^{t}$ for attribute selection, adapting attribute embedding, and refining attribute embedding. As shown in Table 5 and Table 4, results show that all three components are essential. The first and second modules are indispensable, as the absence of either module leads to the model’s inability to perform effective detection. The inclusion of the refining module enhances the model’s detection results. Specifically, although in certain cases, methods that do not involve refinement may result in a slightly lower $FPP$, the detection metrics for each task, such as $AP$, $AP_{.5}$ and $AP_{.75}$, will show significant discrepancies compared to methods that do include refinement, and these discrepancies will continue to widen over time. Therefore, refinement remains necessary.
Table 5: Ablation experiments results of different components on COCO 2017 in two-phase setting 70+10, with the best results highlighted in bold. The “—” indicates that the model lacks object detection capability from the first task, making it meaningless to calculate the FPP value in context.
Sel Ada Ada+Sha Ref Ref+Sha 70 70+10
All All Old $FPP$
$\checkmark$ 0.18 0.17 0.18
$\checkmark$ $\checkmark$ 62.99 0.15 0.16 62.83
$\checkmark$ $\checkmark$ 0.13 0.15 0.16
$\checkmark$ $\checkmark$ $\checkmark$ 63.43 6.75 0.11 63.32
$\checkmark$ $\checkmark$ 62.99 59.71 62.42 0.57
$\checkmark$ $\checkmark$ 0.13 0.12 0.12
$\checkmark$ $\checkmark$ $\checkmark$ 63.43 61.00 63.44 -0.01
On the other hand, we conduct ablation experiments focused on incremental learning. Specifically, we experimented with not sharing attribute information across the three modules and not applying the loss function that enforces the consistency of attribute embedding between the current phase and the previous phase. The results indicate that sharing attribute information during training and applying the loss function that controls phase-to-phase consistency lead to the best performance in incremental object detection.
4.6 Effect of Hyperparameters
The hyperparameter $\lambda_{1}$ serves as a regularization term for the assignment matrix. In our experiments, we fixed $\lambda_{1}$ at 0.01. We conducted extensive experiments on the COCO dataset using the 70+10 setting to analyze the effect of varying two hyperparameters $\lambda_{2}$ and $\lambda_{3}$, which control the consistency of the attribute embedding $\hat{E}_{a}^{t}[\ :Q,:]$ and $\hat{E}_{a}^{t-1}$ respectively. The results, as shown in the figure 3, demonstrate the variation in $AP_{.5}$ values when $\lambda_{2}$ and $\lambda_{3}$ are set to 0.5, 1, 2, and 4. We observed that the change in $AP_{.5}$ is less than 0.02%, indicating minimal impact. Compared to other incremental object detection methods, our approach achieves state-of-the-art performance.
Figure 3: Impact of the hyperparameters $\lambda_{2}$ and $\lambda_{3}$
4.7 Visualizations
The visualization results of incremental object detection are shown in the figure 4. The figure presents the visualization results under the 40+40 experimental setup on the COCO dataset. The red boxes represent the object classes learned by the model in the previous phase, while the green boxes indicate the object classes learned by the model in the current phase. From the visualization results, it can be observed that our method, CASA, demonstrates excellent detection performance for both the current and previously learned classes, effectively achieving incremental learning and overcoming background shift.
Figure 4: Visualization results for 40+40 setting on COCO 2017. The red boxes show object classes learned in the previous phase, while the green boxes represent those learned in the current phase.
5 Conclusions
In this paper, we propose a novel method for incremental object detection (IOD), i.e., CASA, which effectively addresses the challenges of background shift by leveraging vision-language foundation models. By constructing a Class-Agnostic Shared Attribute base, CASA captures and retains common semantic attributes across incremental classes, ensuring consistent generalization across tasks. Our method preserves the original parameters of the pre-trained OWL-ViT model while incorporating parameter-efficient fine-tuning, which only adds 0.7% of parameter storage. This minimal increase in storage significantly enhances the scalability and adaptability of IOD. Extensive experiments conducted on the COCO dataset demonstrate that CASA achieves state-of-the-art performance in both two-phase and multi-phase incremental learning scenarios.